OpenStudy (idku):

A mop is pushed across the floor with a force FF of 50.0 N, at an angle of θ = 50.0°. The mass of the mop head is 3.75 kg. Calculate the acceleration of the mop head if the coefficient of kinetic friction between the head and the floor is μk= 0.400.

OpenStudy (idku):

The image of the situation: http://i.imgur.com/ctpvhx6.jpg

OpenStudy (idku):

Am I doing it correct? The sum of forces in the x: \(\color{#000000}{ \large \displaystyle F_{k ,~x}+F_{{\rm push} ,~x} =ma_x }\) Friction is proportional to normal force: \(\color{#000000}{ \large \displaystyle F_{k}=\mu_kN= \mu_kmg }\) So, the friction-force in the x would be: \(\color{#000000}{ \large \displaystyle F_{k,~x}= \mu_kmg\cos\theta }\) And the force of push in the x is: \(\color{#000000}{ \large \displaystyle F_{{\rm push} ,~x}= F_{{\rm push} } \cos \theta }\) Therefore: (the force of friction is negative, since it counteracts the relative motion. That is why I got the minus there.) \(\color{#000000}{ \large \displaystyle -\mu_kmg\cos\theta + F_{{\rm push} } \cos \theta =ma_x }\) With these particular numbers I have: \(\color{#000000}{ \large \displaystyle -(0.4)(3.75)(9.80)\cos(50) + 50 \cos(50) =3.75~a_x }\) So, the acceleration in the x (and the only acceleration, since the object is obviously not accelerating vertically) is: \(\color{#000000}{ \large \displaystyle a_x=a=\frac{50 \cos(50)-(0.4)(3.75)(9.80)\cos(50)}{3.75} \approx 6.05077 }\)

jhonyy9 (jhonyy9):

so i think how i see your calculi that you are right - good work bye

OpenStudy (idku):

It seemed reasonable to me too, but this answer turned out to be incorrect-:(

jhonyy9 (jhonyy9):

why what will be the right answer ?

OpenStudy (idku):

I don't know. That's what I am trying to find out. I took all of the forces that act in the x-direction into account, but for some reason this answer is incorrect.

OpenStudy (idku):

@Michele_Laino

OpenStudy (michele_laino):

the friction force \(R\) is: \[R = \mu \left( {mg + F\sin \theta } \right)\] so the requested acceleration, is: \[a = \frac{{F\cos \theta - R}}{m}\]

OpenStudy (idku):

\(\color{#000000}{ \large \displaystyle R= \mu_k(mg + F \sin \theta) }\) \(\color{#000000}{ \large \displaystyle a=\frac{F \cos \theta-\mu_k(mg + F \sin \theta) }{m} }\) \(\color{#000000}{ \large \displaystyle a=\frac{50 \cos 50-0.4(3.75\cdot 9.80 + 50 \sin \theta) }{3.75} }\) Like this?

OpenStudy (idku):

Oh, forgot to plug in sin(50)...

OpenStudy (michele_laino):

yes!

OpenStudy (idku):

thanks, I will try that. One question, tho, how did you come up with this formula for R?

OpenStudy (michele_laino):

since the magnitude of R is proportional to the orthogonal pressure exerted by the mop

OpenStudy (michele_laino):

orthogonal with respect to the horizontal Surface, of course

OpenStudy (idku):

Well, or, in two dimensions just perpendicular. So, that would mean sine, not cosine. And proportion is the friction coefficient.

OpenStudy (michele_laino):

yes!

OpenStudy (idku):

That starts to kind of making sense to me. Thanks!

OpenStudy (michele_laino):

:)